5) A block-spring system oscillates in a simple harmonic motion on a frictionless horizontal table. Its displacement varies with time according to x (t) = 0.2 cos (2t - n / 4). The earliest time the particle reaches position x = 0.1 m is


1
Expert's answer
2021-04-13T06:39:55-0400

Explanations & Calculations


  • This form of the equation exhibits that the time is measured with respect to a point located between the center & the amplitude.
  • As usual the distance "\small x" is measured with respect to the center.
  • Then substituting the given value of x in the above equation, the time taken to reach that point can be calculated.

"\qquad\qquad
\begin{aligned}
\small 0.1&=\small 0.2\cos(2t-\frac{\pi }{4})\\
\small \frac{1}{2}=\cos\frac{\pi}{3}&=\small \cos(2t-\frac{\pi}{4})\\
\small \frac{\pi}{3}&=\small 2t-\frac{\pi}{4}\\
\small 2t&=\small \frac{7\pi}{12}\\
\small t&=\small \frac{7\pi}{24}\\
&=\small \bold{1.833\,s}
\end{aligned}" "\qquad\qquad
\begin{aligned}
\text{Extra }&\text{ Information}\\
\small x=0&\to t=\frac{T}{4}\\
\small \frac{\pi}{2}&=\small \Bigg(2\Big(\frac{T}{4}\Big)-\frac{\pi}{4}\Bigg)\\
\small \frac{T}{4}&=\small \frac{3\pi}{8}\approx 1.178\,s\\
\small T&=\small 4.712\,s
\end{aligned}"

  • Extra information,
  • On comparison with the general form of the equation, that is "x(t)=A\cos(\omega t\pm\theta)" , A=0.2 & "\omega =2". Then the periodic time can be found to be,

"\qquad\qquad
\begin{aligned}
\small \omega&=\small \frac{2\pi}{T}=2\\
\small T&=\small \pi \approx3.142\,s 
\end{aligned}"